# A Finite-Time Analysis of Distributed Q-Learning

### v1

@@ 2.3 Switched system
In this paper, we consider a particular system, called the switched affine system [Liberzon, 2005],
@@
3 Error analysis of distributed Q-learning : i.i.d. observation model
In this section, we first consider i.i.d. observation model, which provides simple and clear intuitive results. In the subsequent section, we will extend the result to the Markovian observation model. By an i.i.d. observation model, we refer to a sequence of trajectory $\{(s_{k},{\bm{a}}_{k},s_{k}^{\prime})\}_{k\geq 0}$ where each $(s_{k},{\bm{a}}_{k},s_{k}^{\prime})$ are an i.i.d. random variables. Suppose that each state-action pair is sampled from a distribution $d\in\Delta^{|{\mathcal{S}}\times{\mathcal{A}}|}$, i.e., ${\mathbb{P}}\left[(s_{k},{\bm{a}}_{k})=(s,{\bm{a}})\right]=d(s,{\bm{a}})$ and $s^{\prime}_{k}\sim{\mathcal{P}}(s_{k},{\bm{a}}_{k},\cdot)$. The pseudo-code of the algorithm is given in Algorithm 1 in the Appendix. We will adopt the following standard assumption in the literature [Qu and Wierman, 2020, Li et al., 2024]:
@@ 3.1 Matrix notations
Let us introduce the following vector and matrix notations used throughout the paper to re-write (2) in matrix notations:
${\bm{D}}_{s}$ $:=\text{diag}(d(1,1),d(1,2),\cdots,d(1,|{\mathcal{A}}|))\in\mathbb{R}^{|{\mathcal{A}}|},\quad{\bm{D}}=\text{diag}({\bm{D}}_{1},{\bm{D}}_{2},\dots,{\bm{D}}_{|{\mathcal{S}}|})\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|},$
where $\text{diag}(\cdot)$ is a diagonal matrix whose diagonal elements correspond to the input vector or matrix, and we will denote $d_{\max}=\max_{s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}}d(s,{\bm{a}})$ and $d_{\min}:=\min_{s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}}d(s,{\bm{a}})$. Furthermore, for $i\in[N]$, $o=(s,{\bm{a}},s^{\prime})\in{\mathcal{S}}\times{\mathcal{A}}\times{\mathcal{S}}$ and ${\bm{Q}}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$, we define
@@
${\bm{\Delta}}^{i}({\bm{Q}}):=$ ${\bm{D}}({\bm{R}}^{i}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}}{\bm{Q}}-{\bm{Q}}).$
For simplicity of the notation, we denote $\bm{\delta}^{i}_{k}:=\bm{\delta}^{i}(o_{k},{\bm{Q}}^{i}_{k})$, ${\bm{\Delta}}^{i}_{k}:={\bm{\Delta}}^{i}({\bm{Q}}^{i}_{k})$, and
$\bar{{\bm{Q}}}_{k}$ $:=\begin{bmatrix}{\bm{Q}}^{1}_{k}\\ {\bm{Q}}^{2}_{k}\\ \vdots\\ {\bm{Q}}^{N}_{k}\end{bmatrix},\quad\bar{{\bm{\Pi}}}^{\bar{{\bm{Q}}}_{k}}:=\begin{bmatrix}{\bm{\Pi}}^{{\bm{Q}}^{1}_{k}}&&\\ &\ddots&\\ &&{\bm{\Pi}}^{{\bm{Q}}^{N}_{k}}\end{bmatrix},\quad\bar{\bm{\epsilon}}_{k}(o_{k},\bar{{\bm{Q}}}_{k}):=\begin{bmatrix}\bm{\delta}^{1}(o_{k},{\bm{Q}}^{1}_{k})-{\bm{\Delta}}^{1}({\bm{Q}}^{1}_{k})\\ \bm{\delta}^{2}(o_{k},{\bm{Q}}^{2}_{k})-{\bm{\Delta}}^{2}({\bm{Q}}^{2}_{k})\\ \vdots\\ \bm{\delta}^{N}(o_{k},{\bm{Q}}^{N}_{k})-{\bm{\Delta}}^{N}({\bm{Q}}^{N}_{k})\end{bmatrix}$ (4)
$\bar{{\bm{P}}}$ $:={\bm{I}}_{N}\otimes{\bm{P}},\quad\bar{{\bm{D}}}:={\bm{I}}_{N}\otimes{\bm{D}},\quad\bar{{\bm{W}}}:={\bm{W}}\otimes{\bm{I}}_{|{\mathcal{S}}||{\mathcal{A}}|},\quad\bar{{\bm{R}}}:=\begin{bmatrix}{\bm{R}}^{1}&{\bm{R}}^{2}&\cdots&{\bm{R}}^{N}\end{bmatrix}^{\top},$
where ${\bm{I}}_{N}$ is a $N\times N$ identity matrix, ${\bm{Q}}^{i}_{k}$ is defined in (2). Furthermore, we denote $\bar{\bm{\epsilon}}_{k}:=\bar{\bm{\epsilon}}_{k}(o_{k},\bar{{\bm{Q}}}_{k})$. With the above set of notations, we can re-write the update in (2) as follows:
@@ 3.2 Distributed Q-learning : Error analysis
In this section, we provide a sketch of the proof to bound the error of distributed Q-learning. Let us first decompose the error $\bar{{\bm{Q}}}_{k}-\bm{1}_{N}\otimes{\bm{Q}}^{*}$ into consensus error and optimality error, i.e.,
@@ 3.3 Analysis of Consensus Error
The proof is given in Appendix D.2. The step-size depends on $\min_{i\in[N]}[{\bm{W}}]_{ii}$, which can be considered as a global information. However, considering the method in Example B.1 in Appendix, which requires only local information to construct ${\bm{W}}$, we have $\min_{i\in[N]}[{\bm{W}}]_{ii}\geq\frac{1}{2}$. Therefore, it should be enough to choose $\alpha\leq\frac{1}{2}$. Furthermore, the step-size in many distributed RL algorithms [Zeng et al., 2022b, Wang et al., 2020, Doan et al., 2021, Sun et al., 2020] depend on $\sigma_{2}({\bm{W}})$, which also can be viewed as a global information. Moreover, we can use an agent-specific step-size, i.e., each agent keeps its own step-size, $\alpha_{i}$. Then, we only require $\alpha_{i}<[{\bm{W}}]_{ii}$, which only uses local information. We believe our proofs can be extended to such cases.
@@
$\bm{\Theta}\bar{{\bm{Q}}}_{k+1}=$ $\bar{{\bm{W}}}\bm{\Theta}\bar{{\bm{Q}}}_{k}+\alpha\bm{\Theta}\bar{{\bm{D}}}\left(\bar{{\bm{R}}}+\gamma{\bm{P}}\bar{{\bm{\Pi}}}^{\bar{{\bm{Q}}}_{k}}\bar{{\bm{Q}}}_{k}-\bar{{\bm{Q}}}_{k}\right)+\alpha\bm{\Theta}\bar{\bm{\epsilon}}_{k}$
$=$ $\prod^{k}_{i=0}\bar{{\bm{W}}}^{i}\bm{\Theta}\bar{{\bm{Q}}}_{0}+\alpha\sum_{j=0}^{k}\bar{{\bm{W}}}^{k-j}\bm{\Theta}\left(\bar{{\bm{D}}}\left(\bar{{\bm{R}}}+\gamma\bar{{\bm{P}}}\bar{{\bm{\Pi}}}^{\bar{{\bm{Q}}}_{j}}\bar{{\bm{Q}}}_{j}-\bar{{\bm{Q}}}_{j}\right)+\bar{\bm{\epsilon}}_{j}\right).$ (7)
The first equality follows from the fact that $\bm{\Theta}\bar{{\bm{W}}}=\bar{{\bm{W}}}\bm{\Theta}$. The second equality results from recursively expanding the terms. Now, we are ready to bound $\bm{\Theta}\bar{{\bm{Q}}}_{k+1}$ using the fact that $\left\|\bar{{\bm{W}}}^{i}\bm{\Theta}\right\|_{2}$ for $i\in{\mathbb{N}}$ will decay at a rate of $\sigma_{2}({\bm{W}})$ from Lemma 3.2, and the boundedness of $\bar{{\bm{Q}}}_{k}$ in Lemma 3.3.
@@
The proof is given in Appendix D.3. As we can expect, the convergence rate of the consensus error depends on the $\sigma_{2}({\bm{W}})$ with a constant error bound proportional to $\alpha$. Furthermore, we note that the above result also holds for the Markovian observation model which is introduced in Section 4.
@@ 3.4 Analysis of Optimality Error
We can see that $\bm{\epsilon}^{\mathrm{avg}}_{k}$ is a stochastic term, and we will bound the error caused by this term using concentration inequalities. The consensus error, ${\bm{E}}_{k}$, can be bounded from Theorem 3.4. However, the affine term, ${\bm{b}}_{{\bm{Q}}^{\mathrm{avg}}_{k}}$, does not admit simple bounds. The approach in Lee and He [2020] provides a method to construct a system without an affine term, making the analysis simpler. In details, we introduce a lower and upper comparison system, denoted as ${\bm{Q}}^{\mathrm{avg},l}_{k}$ and ${\bm{Q}}^{\mathrm{avg},u}_{k}$, respectively such that
@@
where the inequality holds in element-wise sense. Letting $\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{k}:={\bm{Q}}^{\mathrm{avg},l}_{k}-{\bm{Q}}^{*}$ and $\tilde{{\bm{Q}}}^{\mathrm{avg}.u}_{k}:={\bm{Q}}^{\mathrm{avg},u}_{k}-{\bm{Q}}^{*}$, a candidate of update that satisfies (11), which is without the affine term ${\bm{b}}_{{\bm{Q}}_{k}}$, is:
@@ 3.5 Final error
The following number of samples are required for $\mathbb{E}\left[\left\|\bar{{\bm{Q}}}_{k}-\bm{1}_{N}\otimes{\bm{Q}}^{*}\right\|_{\infty}\right]\leq\epsilon$:
@@
$\tilde{{\mathcal{O}}}\left(\min\left\{\frac{1}{{\epsilon}^{2}}\frac{d_{\max}^{2}}{(1-\gamma)^{6}d_{\min}^{4}},\frac{1}{{\epsilon}}\frac{d_{\max}^{2}\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\gamma)^{4}d_{\min}^{3}(1-\sigma_{2}({\bm{W}}))}\right\}\right).$
@@
The proof is given in Appendix Section F.3. As the known sample complexity of (single-agent) Q-learning, our bound depends on the factors, $d_{\min}$ and $\frac{1}{1-\gamma}$. The result is improvabale in sense that the known tight dependency for single-agent case is $\frac{1}{(1-\gamma)^{4}d_{\min}}$ by Li et al. [2020]. Furthermore, we note that the dependency on the spectral property of the graph, $\frac{1}{{\epsilon}}\frac{1}{1-\sigma_{2}({\bm{W}})}$ is common in the literature of distributed learning as can be seen in Table 1.
4 Error analysis of distributed Q-learning : Markovian observation model
@@
Now, we consider a Markovian observation model instead of the i.i.d. model. Starting from an initial distribution ${\bm{\mu}}_{0}\in\Delta^{|{\mathcal{S}}||{\mathcal{A}}|}$, the samples are observed from a behavior policy $\beta:{\mathcal{S}}\to\Delta^{|{\mathcal{A}}|}$, i.e., from $(s_{k},{\bm{a}}_{k})$, transition occurs to $s_{k+1}\sim{\mathcal{P}}(s_{k},{\bm{a}}_{k},\cdot)$ and the action is selected by ${\bm{a}}_{k+1}\sim\beta(\cdot\mid s_{k+1})$. This setting is closer to practical scenarios, but poses significant challenges in the analysis due to the dependence between the past observations and current estimates. To overcome this difficulty, we consider the so-called uniformly ergodic Markov chain [Paulin, 2015], which ensures that the Markov chain converges to its unique stationary distribution, ${\bm{\mu}}_{\infty}\in\Delta^{|{\mathcal{S}}||{\mathcal{A}}|}$, exponentially fast in sense of total variation distance, which is defined as $d_{\text{TV}}({\bm{p}},{\bm{q}}):=\frac{1}{2}\sum_{x\in{\mathcal{S}}\times{\mathcal{A}}}|[{\bm{p}}]_{x}-[{\bm{q}}]_{x}|$ where ${\bm{p}},{\bm{q}}\in\Delta^{|{\mathcal{S}}||{\mathcal{A}}|}$. That is, there exist positive real numbers $m,\rho\in\mathbb{R}$ such that for $k\in{\mathbb{N}}$, we have $\max_{s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}}d_{\mathrm{TV}}({\bm{\mu}}_{k}^{s,{\bm{a}}},{\bm{\mu}}_{\infty})\leq m\rho^{k}$, where ${\bm{\mu}}_{k}^{s,{\bm{a}}}:=(({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{P}}^{k}_{\beta})^{\top}$ is the probability distribution of state-action pair after $k$ number of transition occurs starting from $s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}$, and ${\bm{P}}_{\beta}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|\times|{\mathcal{S}}||{\mathcal{A}}|}$ is the transition matrix induced by behavior policy $\beta$, i.e., $({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{P}}_{\beta}({\bm{e}}_{s^{\prime}}\otimes{\bm{e}}_{{\bm{a}}^{\prime}})^{\top}=({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{P}}{\bm{e}}_{s^{\prime}}\cdot\beta({\bm{a}}^{\prime}\mid s^{\prime})$ for $(s,{\bm{a}}),(s^{\prime},{\bm{a}}^{\prime})\in{\mathcal{S}}\times{\mathcal{A}}$. Moreover, we will denote
@@
for ${\epsilon}>0$, and $\tau$ is the so-called mixing time. The concept of mixing time is widely used in the literature [Zeng et al., 2022b, Bhandari et al., 2018]. Note that $\tau$ is approximately proportional to $\log\left(\frac{1}{\alpha}\right)$ [Paulin, 2015] which is provided in Lemma C.7 in the Appendix. This contributes only logarithmic factor to the final error bound. Furthermore, we will denote
@@
Now, we provide the technical difference with the proof of i.i.d. case in Section 3. The challenge in the analysis lies in the fact that $\mathbb{E}\left[\bm{\epsilon}^{\mathrm{avg}}_{k}\middle|\{(s_{t},{\bm{a}}_{t})\}_{t=0}^{k},\bar{{\bm{Q}}}_{0}\right]\neq\bm{0}$ due to Markovian observation scheme. Therefore, we cannot use Azuma-Hoeffding inequality as in the proof of i.i.d. case in the Appendix F.1. Instead, we consider the shifted sequence as in Qu and Wierman [2020]. By shifted sequence, it means to consider the error by the stochastic observation at $k$ with $\bar{{\bm{Q}}}_{k-\tau}$ instead of $\bar{{\bm{Q}}}_{k}$, i.e., ${\bm{w}}_{k,1}:=\bm{\delta}^{\mathrm{avg}}(o_{k},\bar{{\bm{Q}}}_{k-\tau})-{\bm{\Delta}}^{\mathrm{avg}}_{k-\tau,k}(\bar{{\bm{Q}}}_{k-\tau})$ where ${\bm{\Delta}}^{\mathrm{avg}}_{k-\tau,k}(\bar{{\bm{Q}}}_{k}):={\bm{D}}^{s_{k-\tau},{\bm{a}}_{k-\tau}}_{\tau}\frac{1}{N}\sum^{N}_{i=1}\left({\bm{R}}^{i}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}^{i}_{k}}{\bm{Q}}^{i}_{k}-{\bm{Q}}^{i}_{k}\right)$. Then, we have $\mathbb{E}\left[{\bm{w}}_{k,1}\middle|\{(s_{t},{\bm{a}}_{t})\}_{t=0}^{k-\tau},\bar{{\bm{Q}}}_{0}\right]=\bm{0}$. Now, we separately calculate the errors induced by $\{{\bm{w}}_{\tau j+l,1}\}_{j\in\{t\in{\mathbb{N}}\mid\tau t+l\leq k\}}$ for each $10\leq l\leq\tau-$ 1, and invoke the Azuma-Hoeffding inequality. Overall details are given in Appendix G, and we have the following result:

### v2

@@ 2.3 Switched system
In this paper, we consider a system, called the switched affine system [Liberzon, 2005],
@@
3 Error Analysis : i.i.d. observation model
In this section, we first consider i.i.d. observation model, which provides simple and clear intuitive results. In the subsequent section, we will extend the result to the Markovian observation model. By an i.i.d. observation model, we refer to a sequence of trajectory $\{(s_{k},{\bm{a}}_{k},s_{k}^{\prime})\}_{k\geq 0}$ where each $(s_{k},{\bm{a}}_{k},s_{k}^{\prime})$ are an i.i.d. random variables. Suppose that each state-action pair is sampled from a distribution $d\in\Delta^{|{\mathcal{S}}\times{\mathcal{A}}|}$, i.e., ${\mathbb{P}}\left[(s_{k},{\bm{a}}_{k})=(s,{\bm{a}})\right]=d(s,{\bm{a}})$ and $s^{\prime}_{k}\sim{\mathcal{P}}(s_{k},{\bm{a}}_{k},\cdot)$. The pseudo-code of the algorithm is given in Algorithm 1 in the Appendix J. We will adopt the following standard assumption in the literature:
@@ 3.1 Matrix notations
[…]
@@
${\bm{\Delta}}^{i}({\bm{Q}}):=$ ${\bm{D}}({\bm{R}}^{i}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}}{\bm{Q}}-{\bm{Q}}),$
which denotes the TD-error and expected TD-error in vector representation. For simplicity of the notation, we denote $\bm{\delta}^{i}_{k}:=\bm{\delta}^{i}(o_{k},{\bm{Q}}^{i}_{k})$, ${\bm{\Delta}}^{i}_{k}:={\bm{\Delta}}^{i}({\bm{Q}}^{i}_{k})$, and
$\bar{{\bm{Q}}}_{k}$ :$=\begin{bmatrix}{\bm{Q}}^{1}_{k}\\ {\bm{Q}}^{2}_{k}\\ \vdots\\ {\bm{Q}}^{N}_{k}\end{bmatrix},\quad\bar{{\bm{\Pi}}}^{\bar{{\bm{Q}}}_{k}}=\begin{bmatrix}{\bm{\Pi}}^{{\bm{Q}}^{1}_{k}}&&\\ &\ddots&\\ &&{\bm{\Pi}}^{{\bm{Q}}^{N}_{k}}\end{bmatrix},\quad\bar{\bm{\epsilon}}_{k}(o_{k},\bar{{\bm{Q}}}_{k})=\begin{bmatrix}\bm{\delta}^{1}(o_{k},{\bm{Q}}^{1}_{k})-{\bm{\Delta}}^{1}({\bm{Q}}^{1}_{k})\\ \bm{\delta}^{2}(o_{k},{\bm{Q}}^{2}_{k})-{\bm{\Delta}}^{2}({\bm{Q}}^{2}_{k})\\ \vdots\\ \bm{\delta}^{N}(o_{k},{\bm{Q}}^{N}_{k})-{\bm{\Delta}}^{N}({\bm{Q}}^{N}_{k})\end{bmatrix},:$ (4)
$\bar{{\bm{P}}}$ :$={\bm{I}}_{N}\otimes{\bm{P}},\quad\bar{{\bm{D}}}={\bm{I}}_{N}\otimes{\bm{D}},\quad\bar{{\bm{W}}}={\bm{W}}\otimes{\bm{I}}_{|{\mathcal{S}}||{\mathcal{A}}|},\quad\bar{{\bm{R}}}=\begin{bmatrix}{\bm{R}}^{1}&{\bm{R}}^{2}&\cdots&{\bm{R}}^{N}\end{bmatrix}^{\top},:$
where ${\bm{I}}_{N}$ is a $N\times N$ identity matrix, ${\bm{Q}}^{i}_{k}$ is defined in (2). Moreover, we denote $\bar{\bm{\epsilon}}_{k}:=\bar{\bm{\epsilon}}_{k}(o_{k},\bar{{\bm{Q}}}_{k})$. With the above set of notations, we can re-write the update in (2) as follows:
@@ 3.2 Distributed Q-learning : Error analysis
In this section, we provide a sketch of the proof to bound the error of distributed Q-learning. Let us first decompose the error $\bar{{\bm{Q}}}_{k}-\bm{1}_{N}\otimes{\bm{Q}}^{*}$ into consensus error and optimality error:
@@ 3.3 Analysis of Consensus Error
The proof is given in Appendix D.2. The step-size depends on $\min_{i\in[N]}[{\bm{W}}]_{ii}$, which can be considered as a global information. However, considering the method in Example B.1 in Appendix, which requires only local information to construct ${\bm{W}}$, we have $\min_{i\in[N]}[{\bm{W}}]_{ii}\geq\frac{1}{2}$. Therefore, it should be enough to choose $\alpha\leq\frac{1}{2}$. Furthermore, the step-size in many distributed RL algorithms [Zeng et al., 2022b, Wang et al., 2020, Doan et al., 2021, Sun et al., 2020] depend on $\sigma_{2}({\bm{W}})$, which also can be viewed as a global information. Moreover, we can use an agent-specific step-size, i.e., each agent keeps its own step-size, $\alpha_{i}$. Then, we only require $\alpha_{i}<[{\bm{W}}]_{ii}$, which only uses local information.
@@
$\bm{\Theta}\bar{{\bm{Q}}}_{k+1}=$ $\prod^{k}_{i=0}\bar{{\bm{W}}}^{i}\bm{\Theta}\bar{{\bm{Q}}}_{0}+\alpha\sum_{j=0}^{k}\bar{{\bm{W}}}^{k-j}\bm{\Theta}\left(\bar{{\bm{D}}}\left(\bar{{\bm{R}}}+\gamma\bar{{\bm{P}}}\bar{{\bm{\Pi}}}^{\bar{{\bm{Q}}}_{j}}\bar{{\bm{Q}}}_{j}-\bar{{\bm{Q}}}_{j}\right)+\bar{\bm{\epsilon}}_{j}\right).$ (7)
The equality results from recursively expanding the terms. Now, we are ready to bound $\bm{\Theta}\bar{{\bm{Q}}}_{k+1}$ using the fact that $\left\|\bar{{\bm{W}}}^{i}\bm{\Theta}\right\|_{2}$ for $i\in{\mathbb{N}}$ will decay at a rate of $\sigma_{2}({\bm{W}})$ from Lemma 3.2, and the boundedness of $\bar{{\bm{Q}}}_{k}$ in Lemma 3.3.
@@
The proof is given in Appendix D.3. As we can expect, the convergence rate of the consensus error depends on the $\sigma_{2}({\bm{W}})$ with a constant error bound proportional to $\alpha$. Furthermore, we note that the above result also holds for the Markovian observation model in Section 4.
@@ 3.4 Analysis of Optimality Error
We can see that $\bm{\epsilon}^{\mathrm{avg}}_{k}$ is a stochastic term, and we will bound the error caused by this term using concentration inequalities. The consensus error, ${\bm{E}}_{k}$, can be bounded from Theorem 3.4. However, the affine term, ${\bm{b}}_{{\bm{Q}}^{\mathrm{avg}}_{k}}$, does not admit simple bounds. The approach in Lee and He [2020] provides a method to construct a system without an affine term, making the analysis simpler. In details, we introduce a lower and upper comparison system, denoted as ${\bm{Q}}^{\mathrm{avg},l}_{k}$ and ${\bm{Q}}^{\mathrm{avg},u}_{k}$, respectively such that the following element-wise inequaltiy holds:
@@
Letting $\tilde{{\bm{Q}}}^{\mathrm{avg},l}_{k}:={\bm{Q}}^{\mathrm{avg},l}_{k}-{\bm{Q}}^{*}$ and $\tilde{{\bm{Q}}}^{\mathrm{avg}.u}_{k}:={\bm{Q}}^{\mathrm{avg},u}_{k}-{\bm{Q}}^{*}$, a candidate of update that satisfies (11), which is without the affine term ${\bm{b}}_{{\bm{Q}}_{k}}$, is:
@@ 3.5 Final error
Suppose $\alpha=\tilde{{\mathcal{O}}}\left(\min\left\{\frac{(1-\gamma)^{5}d_{\min}^{3}}{R_{\max}^{2}d_{\max}^{2}}{\epsilon}^{2},\frac{(1-\gamma)^{3}d_{\min}^{2}(1-\sigma_{2}({\bm{W}}))}{R_{\max}d_{\max}^{2}\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{\epsilon}\right\}\right)$. Then, the following number of samples are required for $\mathbb{E}\left[\left\|\bar{{\bm{Q}}}_{k}-\bm{1}_{N}\otimes{\bm{Q}}^{*}\right\|_{\infty}\right]\leq\epsilon$:
@@
$\tilde{{\mathcal{O}}}\left(\max\left\{\frac{1}{{\epsilon}^{2}}\frac{d_{\max}^{2}}{(1-\gamma)^{6}d_{\min}^{4}},\frac{1}{{\epsilon}}\frac{d_{\max}^{2}\sqrt{|{\mathcal{S}}||{\mathcal{A}}|}}{(1-\gamma)^{4}d_{\min}^{3}(1-\sigma_{2}({\bm{W}}))}\right\}\right).$
@@
The proof is given in Appendix Section F.3. As the known sample complexity of (single-agent) Q-learning, our bound depends on the factors, $d_{\min}$ and $\frac{1}{1-\gamma}$. The result is improvable in sense that the known tight dependency for single-agent case is $\frac{1}{(1-\gamma)^{4}d_{\min}}$ by Li et al. [2020]. Furthermore, we note that the dependency on the spectral property of the graph, $\frac{1}{{\epsilon}}\frac{1}{1-\sigma_{2}({\bm{W}})}$ is common in the literature of distributed learning as can be seen in Table 1.
4 Error Analysis : Markovian observation model
@@
Now, we consider a Markovian observation model instead of the i.i.d. model. Starting from an initial distribution ${\bm{\mu}}_{0}\in\Delta^{|{\mathcal{S}}||{\mathcal{A}}|}$, the samples are observed from a behavior policy $\beta:{\mathcal{S}}\to\Delta^{|{\mathcal{A}}|}$, i.e., from $(s_{k},{\bm{a}}_{k})$, transition occurs to $s_{k+1}\sim{\mathcal{P}}(s_{k},{\bm{a}}_{k},\cdot)$ and the action is selected by ${\bm{a}}_{k+1}\sim\beta(\cdot\mid s_{k+1})$. This setting is closer to practical scenarios, but poses significant challenges in the analysis due to the dependence between the past observations and current estimates. To overcome this difficulty, we consider the so-called uniformly ergodic Markov chain [Paulin, 2015], which ensures that the Markov chain converges to its unique stationary distribution, ${\bm{\mu}}_{\infty}\in\Delta^{|{\mathcal{S}}||{\mathcal{A}}|}$, exponentially fast in sense of total variation distance, which is defined as $d_{\text{TV}}({\bm{p}},{\bm{q}}):=\frac{1}{2}\sum_{x\in{\mathcal{S}}\times{\mathcal{A}}}|[{\bm{p}}]_{x}-[{\bm{q}}]_{x}|$ where ${\bm{p}},{\bm{q}}\in\Delta^{|{\mathcal{S}}||{\mathcal{A}}|}$. That is, there exist positive real numbers $m\in\mathbb{R}$ and $\rho\in(0,1)$ such that we have $\max_{s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}}d_{\mathrm{TV}}({\bm{\mu}}_{k}^{s,{\bm{a}}},{\bm{\mu}}_{\infty})\leq m\rho^{k}$, where ${\bm{\mu}}_{k}^{s,{\bm{a}}}:=(({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{P}}^{k}_{\beta})^{\top}$ is the probability distribution of state-action pair after $k$ number of transition occurs starting from $s,{\bm{a}}\in{\mathcal{S}}\times{\mathcal{A}}$, and ${\bm{P}}_{\beta}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|\times|{\mathcal{S}}||{\mathcal{A}}|}$ is the transition matrix induced by behavior policy $\beta$, i.e., $({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{P}}_{\beta}({\bm{e}}_{s^{\prime}}\otimes{\bm{e}}_{{\bm{a}}^{\prime}})^{\top}=({\bm{e}}_{s}\otimes{\bm{e}}_{{\bm{a}}})^{\top}{\bm{P}}{\bm{e}}_{s^{\prime}}\cdot\beta({\bm{a}}^{\prime}\mid s^{\prime})$. Moreover, we will denote
@@
for ${\epsilon}>0$, and $\tau$ is the so-called mixing time. The concept of mixing time is widely used in the literature [Zeng et al., 2022b, Bhandari et al., 2018]. Note that $\tau$ is approximately proportional to $\log\left(\frac{1}{\alpha}\right)$, which is provided in Lemma C.7 in the Appendix. This contributes only logarithmic factor to the final error bound. Furthermore, we will denote
@@
Now, we provide the technical difference with the proof of i.i.d. case in Section 3. The challenge in the analysis lies in the fact that $\mathbb{E}\left[\bm{\epsilon}^{\mathrm{avg}}_{k}\middle|\{(s_{t},{\bm{a}}_{t})\}_{t=0}^{k},\bar{{\bm{Q}}}_{0}\right]\neq\bm{0}$ due to Markovian observation scheme. Therefore, we cannot use Azuma-Hoeffding inequality as in the proof of i.i.d. case in the Appendix F.1. Instead, we consider the shifted sequence as in Qu and Wierman [2020]. By shifted sequence, it means to consider the error by the stochastic observation at $k$ with $\bar{{\bm{Q}}}_{k-\tau}$ instead of $\bar{{\bm{Q}}}_{k}$, i.e., ${\bm{w}}_{k,1}:=\bm{\delta}^{\mathrm{avg}}(o_{k},\bar{{\bm{Q}}}_{k-\tau})-{\bm{\Delta}}^{\mathrm{avg}}_{k-\tau,k}(\bar{{\bm{Q}}}_{k-\tau})$ where ${\bm{\Delta}}^{\mathrm{avg}}_{k-\tau,k}(\bar{{\bm{Q}}}_{k}):={\bm{D}}^{s_{k-\tau},{\bm{a}}_{k-\tau}}_{\tau}\frac{1}{N}\sum^{N}_{i=1}\left({\bm{R}}^{i}+\gamma{\bm{P}}{\bm{\Pi}}^{{\bm{Q}}^{i}_{k}}{\bm{Q}}^{i}_{k}-{\bm{Q}}^{i}_{k}\right)$. Then, we have $\mathbb{E}\left[{\bm{w}}_{k,1}\middle|\{(s_{t},{\bm{a}}_{t})\}_{t=0}^{k-\tau},\bar{{\bm{Q}}}_{0}\right]=\bm{0}$. Now, we separately calculate the errors induced by $\{{\bm{w}}_{\tau j+l,1}\}_{j\in\{t\in{\mathbb{N}}\mid\tau t+l\leq k\}}$ for each $0\leq l\leq\tau-1$, and invoke the Azuma-Hoeffding inequality. Overall details are given in Appendix G, and we have the following result: